OpenStudy (anonymous):

hi i need help to work out this problem, thanks in advance. let p(z)=f(z-1) and f(z)=z^3-8 express p(z) in the form p(z)=z^3+az^2+bz+c hence solve p(z)=0

OpenStudy (anonymous):

p(z) = f(z-1) therefore: p(z) = (z-1)^3 - (z-1) this expands to: p(z) = (z^2 - 2z + 1)(z-1) - z + 1 = z^3 - z^2 - 2z^2 + 2z + z - 1 - z + 1 = z^3 - 3z^2 + 2z Now, can you solve the cubic equation: z^3 - 3z^2 + 2z = 0

OpenStudy (anonymous):

ok so how do work out the cubic equation?

OpenStudy (anonymous):

Factor it.

OpenStudy (anonymous):

could you please show me how?

OpenStudy (anonymous):

You know that the factorization has to be of form: (x + a)(x + b)(x + c) I recommend this guide: http://www.wikihow.com/Factor-a-Cubic-Polynomial

OpenStudy (anonymous):

Yeah. So, to get the solutions for z, just solve the equations: z - 2 = 0 z - 1 = 0 z = 0

OpenStudy (anonymous):

which the equals 2,1,0 is that correct

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

I have a lifetime.

OpenStudy (anonymous):

Is there a picture and/or more information?

OpenStudy (anonymous):

Is there a definition for OA, OB, OC in the standard basis?

OpenStudy (anonymous):

Yeah. So I'm assuming it's defining OA = <2, 0>, OB = <1, 0>, OC = <0, 0>?

OpenStudy (anonymous):

yeah that seems right and relevant to the rest of the question

OpenStudy (anonymous):

Do you know how to get the polar forms of those vectors?

OpenStudy (anonymous):

Do you think I did not explain the previous question properly? If so, I apologize.

OpenStudy (anonymous):

no ,not at all you did great but just that question about vectors it would be good if you explained it thanks

OpenStudy (anonymous):

Okay, so you have a vector in the rectangular form. Let's take <2, 0>, OA. Obviously, we know immediately that the angle must be 0. We can check this with: $$\tan^{-1}(0)$$ Normally we would used atan2, but it's fine. We're taking the inverse tangent of 0, the y-coordinate.

OpenStudy (anonymous):

The arctangent of the y-coord gives us the angle, theta. Now to find the magnitude we just use \(\sqrt{x^2 + y^2}\)

OpenStudy (anonymous):

sqrt(2^2 + 0^2) = sqrt(4) = 2 So the polar form of the vector is (2, 0 degrees), because the magnitude is 2 and theta = 0

OpenStudy (anonymous):

Alright. Can you think of a way to express this in terms of \(\vec i, \vec j\)?